I have been given directions to "Solve the equation for the given variable"
my problem is ' P=2l+2w for w '. I tried subtracting 2l from both expressions, but i come up with the wrong answer. The answer is suppose to be ' w=1/2p-l ', but i do not know how to get that answer.

Answer by Flannery(124)

after subtracting 2l divide by 2 to isolate the w.
p=2l+2w
after subtracting 2l from both sides
p-2l=2w
after dividing both sides by 2
p-2l/2=w
it's really 1p/2-2l/2=w
there are two whole fractions here, seperated by a minus sign.
they are 1p/2 and 2l/2
in the second term the 2/2=1 and dosn't have to be written out (the 2's cancel)
so its 1p/2-l=w
or, you could express it is 1/2p-l=w, which is the same as saying w=1/2p-l
w=p/2-l is also correct.